OpenStudy (oswaldmurphy):

There are 20 parrots at the animal sanctuary. Their population is increasing at a rate of 15% per year. There are also 24 snakes at the sanctuary. Each year 4 more snakes are born. Part A: Write functions to represent the numbers of parrots and snakes at the sanctuary throughout the years. (4 points) Part B: How many parrots are at the sanctuary after 10 years? How many snakes are at the sanctuary after the same number of years? Assume there are no deaths to the animals during this time. (2 points)

OpenStudy (oswaldmurphy):

Part C: After approximately how many years is the number of parrots and snakes the same? Justify your answer mathematically. (4 points)

OpenStudy (oswaldmurphy):

@zepdrix @sammixboo @pooja195

OpenStudy (oswaldmurphy):

would the first be, parrots: f(x) = 20(1.15)^x and Snakes: f(x) = 24(4)^x?

OpenStudy (welshfella):

the growth in population of the parrots is geometric yes that is population of parrots in year x

OpenStudy (oswaldmurphy):

So, Part A: parrots: f(x) = 20(1.15)^x and Snakes: f(x) = 24(4)^x

OpenStudy (oswaldmurphy):

@mathstudent55

OpenStudy (oswaldmurphy):

No snakes would be 24 + 4x..

OpenStudy (oswaldmurphy):

B would be 80 and 64, C you would just solve for x by making the equations = each other.

OpenStudy (welshfella):

sorry i got called away

OpenStudy (agent0smith):

Note that for C you can only solve by graphing, not by hand.

OpenStudy (welshfella):

yes you can use desmos graphing calculator online. draw 2 graphs y = 24+ 4x and y = 20(1.15)^x The required year will be where they intersect.